find the equation of the line that passes through (2, -3) and parallel to 3x + y = 8
Answer by jim_thompson5910(35256)

Anything parallel to 3x + y = 8 will be of the form 3x + y = C.


3x + y = C


3(2) + (-3) = C ... plug in (x,y) = (2,-3)


3(2) - 3 = C


6 - 3 = C


3 = C


C = 3


So the equation of the parallel line is 3x + y = 3